OpenStudy (anonymous):

8 times the square of a nonzero number is equal to 96 times the number.

OpenStudy (anonymous):

8 * x^2 = 96x

OpenStudy (anonymous):

There are two solutions, x = 0 and x = 12. 8 * 0^2 = 96(2) 0 = 0 8 * 12^2 = 96(12) 1152 = 1152

OpenStudy (anonymous):

Because the question asks for a non zero number, the answer is 12.